# Bar formation and evolution in the cosmological context: Inputs from the Auriga simulations

### v2

@@ Abstract
Galactic bars drive the internal evolution of spiral galaxies, while their formation is tightly coupled to the properties of their host galaxy and dark matter halo. To explore what drives bar formation in the cosmological context and how these structures evolve throughout cosmic history, we use the Auriga suite of magneto-hydrodynamical cosmological zoom-in simulations. We find that bars are robust and long-lived structures, and we recover a decreasing bar fraction with increasing redshift which plateaus around $\sim 20\%$ at $z\sim 3$. We find that bars which form at low and intermediate redshifts grow longer with time, while bars that form at high redshifts are born ‘saturated’ in length, likely due to their merger-induced formation pathway. This leads to a larger bar-to-disc size ratio at high redshifts as compared to the local Universe. We subsequently examine the multi-dimensional parameter space thought to drive bar formation. We find that barred galaxies tend to have lower Toomre $Q$ values at the time of their formation, while we do not find a difference in the gas fraction of barred and unbarred populations when controlling for stellar mass. Barred galaxies tend to be more baryon-dominated at all redshifts, assembling their stellar mass earlier, while galaxies that are baryon-dominated but that do not host a bar, have a higher ex-situ bulge fraction. We explore the implications of the baryon-dominance of barred galaxies on the Tully-Fisher relation, finding an offset from the unbarred relation; confirming this in observations would serve as additional evidence for dark matter, as this behaviour is not readily explained in modified gravity scenarios.
@@ 1 Introduction
The aforementioned studies have largely been carried out using isolated N-body simulations. These lend themselves to exploring the internal evolution of galaxies, due to the higher resolution they offer, while they also enable one to isolate specific physical properties to study. However, it is becoming increasingly clear that the mechanisms responsible for bar formation and evolution cannot be decoupled from the cosmological context, since processes such as mergers (e.g. Athanassoula et al. 2017), interactions (e.g. Łokas 2019), external gas accretion (e.g. Bournaud & Combes 2002), the physics of star formation and various feedback processes (e.g. Irodotou et al. 2022), as well as the assembly and properties of the dark matter halos themselves (e.g. Collier et al. 2018), all play a role in the formation and evolution of bars. Recent advances in terms of numerical resolution and modeling of the physics of baryons have made it possible to begin to use cosmological simulations to explore the internal dynamical evolution of galaxies both in ‘zoom-in simulations’ (e.g. Scannapieco & Athanassoula 2012; Kraljic et al. 2012; Buck et al. 2018; Gargiulo et al. 2019; Fragkoudi et al. 2020, 2021; Merrow et al. 2023) as well as in ‘large box’ cosmological simulations (e.g. Algorry et al. 2017; Peschken & Łokas 2019; Rosas-Guevara et al. 2020; Roshan et al. 2021; Frankel et al. 2022; Gargiulo et al. 2022; Izquierdo-Villalba et al. 2022; López et al. 2024). It is worth noting that all these simulations, which use different prescriptions for modelling the baryonic physics of galaxy formation and evolution, often lead to substantially different predictions for the fraction of bars (e.g. finding very few bars at low redshifts $z\sim 0.25$; Reddish et al. 2022) as well as for their intrinsic properties (e.g. weak and transient bars; Ansar et al. 2023). As the galaxy formation (or ‘baryon’ physics) models become more advanced, we can use the properties of bars to constrain these models, since bars act as a litmus test of whether the dynamics of observed discs are being adequately modelled in cosmological simulations.
@@
Figure 1: Face-on RGB stellar light projections of the barred galaxies in the Auriga simulations. The panels have a size of 50 kpc×$\times$50 kpc.
@@ 2 The Auriga Simulations
Figure 3: The $m=2$ Fourier mode of the surface density as a function of lookback time for galaxies which are barred at $z=0$, colour-coded by the formation time of the bar (dark purple: bar forms at larger lookback time, light yellow: bar forms close to $z=0$). The galaxies that end up barred at $z=0$ develop their bars at various times in the simulation, some forming early on and staying barred, while others have bars that develop only at late times.
@@
Figure 5: Top: Bar strength at $z=0$ versus the age of the bar. The colour-coding gives the bar strength (used in subsequent plots), and the size of the points indicates the length of the bar at $z=0$. We indicate the Pearson coefficient $r$ and the p-value in the figure. Bottom: Bar formation time as a function of the time at which the galaxy becomes baryon-dominated. The dashed line indicates is 1-to-1 line.
Figure 6: Bar formation time as a function of various parameters: Left: Bar formation time as a function of the stellar-to-dark matter ratio inside 5 kpc at $z=0$. The numbers indicate the Pearson coefficient $r$ and the p-value. Middle: Bar formation time as a function of the stellar mass of the galaxy at $z=0$. Right: Bar formation time as a function of the stellar mass of the galaxy at the time of bar formation. There is a mild correlation between the bar formation time and the stellar-to-dark matter ratio at $z=0$, while there is no correlation between bar formation time and stellar mass at $z=0$ (i.e no obvious signs of mass downsizing within this mass range). There is a trend of decreasing stellar mass for earlier bar formation time, as galaxies at higher redshifts have lower stellar masses.
@@
Figure 9: The bar length-to-disc size ratio as a function of the formation time of the bar. Left: The ratio at the time of bar formation; Middle: The ratio at $z=0$; Right: The change in the ratio from bar formation time to $z=0$. We see that the ratio of bar length-to-disc size is larger at the time of bar formation for early forming bars, and decreases at $z=0$ due to the increasing disc size.
@@ 2.2 Analysis
We note that for estimating the bar fraction in the analysis that follows, we also visually inspect the galaxies, to ensure that if $A_{2}\geq 0.25$, this is not due, for example, to a transient off-centering of the disc due to a merger, which can occur frequently at high redshifts.
@@
Bar length, $R_{\rm bar}$: There are various methods for estimating the length of the bar, from ellipse-fitting methods used widely in observational studies, to methods using the $m=2$ Fourier mode of the surface density, which are more commonly employed in numerical studies (see e.g. Erwin 2005; Gadotti 2011; Hilmi et al. 2020; Fragkoudi et al. 2021; Petersen et al. 2023). Here we use a method which is well-suited for estimating the bar length in an automated way for simulated galaxies, across a large number of snapshots, i.e. we use the radius at which $A_{2}$ falls below 70% of its maximum value. We note that we have tested various methods for obtaining the bar length, such as a definition based on the constancy of the phase of the $m=2$ Fourier mode. We found this latter method to be more prone to noisy bar length estimates. This lead us to employ the former method, as it is more robust (see also Figure B2 of Fragkoudi et al. 2021).
@@ 3.1 Global Properties of Barred Galaxies
Mass Downsizing and Correlations with $t_{\rm bf}$: In the top panel of Fig. 5 we show the bar strength at $z=0$ as a function of the bar formation time. We find a weak correlation between the bar strength at $z=0$ and the age (or formation time) of the bar, indicating that older bars tend to be slightly stronger. In the bottom panel of the Figure we show the bar formation time as a function of the time at which the galaxy becomes baryon dominated (see Section 2.2 for details). We see that there is a correlation between these quantities, suggesting that galaxies form a bar when the galaxy passes a threshold of baryon-dominance (and see also Section 4.1).
@@ 3.2 Evolution of Bar Properties
In Figure 7 we show the evolution of bar length as a function of time. In particular, we group together bars according to their formation time, i.e. we select bars that form between lookback times of $t_{\rm lb}=0-2.7$ Gyr ago (dark purple), $t_{\rm lb}=2.8-4.9$ Gyr (dark blue), $t_{\rm lb}=5-7.1$ Gyr (dark green) and $t_{\rm lb}=7.2-11$ Gyr (light green), as indicated by the figure inset. These bins are chosen such that they split the sample into 4 groups with approximately the same number of galaxies. We find that bars that form at low redshifts, tend to form shorter bars which subsequently grow longer over time. On the other hand, bars that form at higher redshifts tend to already be quite long when they form, and their length does not evolve much over the galaxy’s lifetime. This is likely due to the fact that galaxies that form bars at high-$z$ are triggered by a merger, and therefore form long and “saturated”, i.e. they are not able to grow by exchanging more angular momentum. Indications of this can be seen in the inset of Fig. 7, where we show the time of one of the three most massive mergers in the galaxy, versus the time of bar formation. We see that the galaxies that form bars at high redshift all do so around the time of one of their most massive mergers.
@@ 4.1 Baryon dominance
Figure 11: Build-up of stellar and dark matter mass for barred and unbarred galaxies. Left: The ratio of stellar over dark matter mass within 5 kpc for barred (red) and unbarred (blue) galaxies in the Auriga suite of simulations. On average, for barred galaxies, the stellar mass dominates over the dark matter mass within the central 5 kpc by $z\sim 0$, which is in contrast to unbarred galaxies which tend to be more dark matter dominated over all cosmic times. Right: The build-up of stellar (solid) and dark matter (dashed) mass within 5 kpc for barred (red) and unbarred (blue) galaxies. Barred galaxies assemble their stellar mass earlier on and form on average 3 Gyr before their unbarred counterpart.
@@
Fig. 11 explores the ratio of stellar-to-dark matter within 5 kpc ($f_{\rm bd}$) as a function of time. In the left panel of the figure we see that already at earlier lookback times, the barred galaxies are more baryon dominated than the unbarred sample. In the right panel we show the growth of the stellar (solid) and dark matter (dashed) mass of the barred (red) and unbarred (blue) galaxies. We see that barred galaxies assemble their stellar mass more rapidly than unbarred galaxies (see also Rosas-Guevara et al. 2020), while the dark matter mass is assembled similarly when comparing the barred and unbarred galaxies. We quantify this by exploring when the samples assemble 50% of their stellar mass, and find that barred galaxies assemble their stellar mass on average $\sim$3 Gyr earlier than unbarred galaxies. Barred galaxies therefore tend to grow their stellar component more rapidly than unbarred galaxies, likely forming an early massive stellar disc, which is conducive to bar formation.
@@ 5 Discussion
Figure 16: The stellar Tully-Fisher relation for galaxies in Auriga barred galaxies (excluding interacting) and unbarred galaxies. Left: The grey dot-dashed line shows the relation given by fitting all galaxies together, while the shaded region indicates the intrinsic vertical scatter $\sigma_{y}$. Right: The red (blue) dashed line shows the relation when fitting only the barred (unbarred) galaxies. The shaded regions indicate the uncertainties from boostrapping (see text).We find that the TF relation for barred galaxies is offset from that of the unbarred galaxies.
Figure 17: The multi-dimensional parameter space of bar formation: dark matter halo mass $M_{\rm DM}$, stellar mass $M_{\star}$, baryon-dominance $f_{\rm bd}$, fraction of accreted stars in the central regions (bulge) $f_{\rm acc}$, the Toomre parameter within the half-mass radius $Q_{r_{50}}$ at the time of bar formation and the gas fraction at $z=0$ $f_{\rm gas}$. The red (blue) points indicate the barred (unbarred) sample, including interacting galaxies. We see that the barred-unbarred populations separate out for the baryon dominance, accreted bulge fraction and Toomre $Q$ parameters.
@@ 5.1 Implications for the Tully-Fisher relation
The Tully-Fisher relation (TF; Tully & Fisher 1977) is an empirical relation between the rotation velocity of galaxies – often the radius at which the rotation curve flattens out – versus their luminosity. This relation has low intrinsic scatter, making it a powerful tool for determining distances to galaxies, while it has also proved useful for testing galaxy formation models in $\Lambda$CDM (Mo et al., 1998) and theories of modified gravity, such as Modified Newtonian Dynamics, or MOND (see e.g. Milgrom 1983; McGaugh 2012). In Fig. 16 we explore the consequences of our findings in Section 4.1 – i.e. the fact that barred galaxies tend to be more baryon-dominated than unbarred galaxies – on the stellar TF relation (STFR; in Appendix Bar formation and evolution in the cosmological context: Inputs from the Auriga simulations we show results for the Baryonic Tully Fisher relation). We obtain the circular velocity of the galaxies in our sample at 30 kpc – which corresponds roughly to the flat part of the rotation curve for the galaxies we are exploring, i.e. $V_{\rm flat}$. We then fit the relation $\log{M_{\star}}=m\log{V_{\rm flat}}+y_{0}$ to our sample of barred galaxies (excluding interracting galaxies) and unbarred galaxies. In the left panel of Fig. 16 we show a least squares fit to this sample (grey dashed curve). The vertical scatter of the relation in our sample is $\sigma_{y}=0.05$dex, as indicated with the grey shaded region. In the right panel of the Figure we fit separately the barred (red) and unbarred (blue) samples. For these fits, we fix the slope of the relation to be that obtained for the total sample and we perform bootstrapping to assess the uncertainties in the fits. The 95% percentile uncertainty in the intercept, as obtained form our fits, is shown with the shaded regions. We find an offset in the STFR of the barred galaxies as compared to that of the unbarred galaxies, of the order of $\Delta y_{0}=0.05$dex. This offset is comparable to the intrinsic scatter of the entire sample, suggesting that, while challenging, it could in principle be observed.

### v3

@@ Abstract
Galactic bars drive the internal evolution of spiral galaxies, while their formation is tightly coupled to the properties of their host galaxy and dark matter halo. To explore what drives bar formation in the cosmological context and how these structures evolve throughout cosmic history, we use the Auriga suite of magneto-hydrodynamical cosmological zoom-in simulations. We find that bars are robust and long-lived structures, and we recover a decreasing bar fraction with increasing redshift which plateaus around $\sim 20\%$ at $z\sim 3$. We find that bars which form at low and intermediate redshifts grow longer with time, while bars that form at high redshifts are born ‘saturated’ in length, likely due to their merger-induced formation pathway. This leads to a larger bar-to-disc size ratio at high redshifts as compared to the local Universe. We subsequently examine the multi-dimensional parameter space thought to drive bar formation. We find that barred galaxies tend to have lower Toomre $Q$ values at the time of their formation, while we do not find a difference in the gas fraction of barred and unbarred populations when controlling for stellar mass. Barred galaxies tend to be more baryon-dominated at all redshifts and assemble their stellar mass earlier, while galaxies that are baryon-dominated but that do not host a bar, have a higher ex-situ bulge fraction. We explore the implications of the baryon-dominance of barred galaxies on the Tully-Fisher relation, finding an offset from the unbarred relation; confirming this in observations would serve as additional evidence for dark matter, as this behaviour is not readily explained in modified gravity scenarios.
@@ 1 Introduction
The aforementioned studies have largely been carried out using isolated N-body simulations. These lend themselves to exploring the internal evolution of galaxies, due to the higher resolution they offer, while they also enable one to isolate specific physical properties to study. However, it is becoming increasingly clear that the mechanisms responsible for bar formation and evolution cannot be decoupled from the cosmological context, since processes such as mergers (e.g. Athanassoula et al. 2017), interactions (e.g. Łokas 2019), external gas accretion (e.g. Bournaud & Combes 2002), the physics of star formation and various feedback processes (e.g. Irodotou et al. 2022), as well as the assembly and properties of the dark matter halos themselves (e.g. Collier et al. 2018), all play a role in the formation and evolution of bars. Recent advances in terms of numerical resolution and modeling of the physics of baryons have made it possible to begin to use cosmological simulations to explore the internal dynamical evolution of galaxies both in ‘zoom-in simulations’ (e.g. Scannapieco & Athanassoula 2012; Kraljic et al. 2012; Buck et al. 2018; Gargiulo et al. 2019; Fragkoudi et al. 2020, 2021; Merrow et al. 2024) as well as in ‘large box’ cosmological simulations (e.g. Algorry et al. 2017; Peschken & Łokas 2019; Rosas-Guevara et al. 2020; Roshan et al. 2021; Frankel et al. 2022; Gargiulo et al. 2022; Izquierdo-Villalba et al. 2022; López et al. 2024). It is worth noting that all these simulations, which use different prescriptions for modelling the baryonic physics of galaxy formation and evolution, often lead to substantially different predictions for the fraction of bars (e.g. finding very few bars at low redshifts $z\sim 0.25$; Reddish et al. 2022) as well as for their intrinsic properties (e.g. weak and transient bars; Ansar et al. 2023). As the galaxy formation (or ‘baryon’ physics) models become more advanced, we can use the properties of bars to constrain these models, since bars act as a litmus test of whether the dynamics of observed discs are being adequately modelled in cosmological simulations.
@@
Figure 1: Face-on RGB stellar light projections of the barred galaxies in the Auriga simulations at z=0. The panels have a size of 50 kpc×$\times$50 kpc.
@@ 2 The Auriga Simulations
Figure 3: The $m=2$ Fourier mode of the surface density as a function of lookback time for galaxies which are barred at $z=0$. The lines are sorted by the formation time of the bar (dark purple: bar forms at larger lookback time, light yellow: bar forms close to $z=0$). The galaxies that end up barred at $z=0$ develop their bars at various times in the simulation, some forming early on and staying barred, while others have bars that develop only at late times.
@@
Figure 5: Top: Bar strength at $z=0$ versus the age of the bar. The colour-coding gives the bar strength (used in subsequent plots). The points outlined in black denote the ‘interacting’ sample (see text and Table 1). Middle: Bar formation time as a function of the time at which the galaxy becomes baryon-dominated. We indicate the Pearson coefficient, $r$, and the p-value for the entire sample of galaxies in the top left. Bottom: Bar formation time as a function of the time at which the galaxy assembles 50% of its stellar mass. The dashed line indicates the linear regression fit for the non-interacting galaxies (see text).
Figure 6: Bar formation time as a function of various parameters. Black circles denote the interacting galaxies and the Pearson coefficient, $r$, and p-value are indicated at the top of each panel: Left: Bar formation time as a function of the stellar-to-dark matter ratio inside 5 kpc at $z=0$. Middle: Bar formation time as a function of the stellar mass of the galaxy at $z=0$. Right: Bar formation time as a function of the stellar mass of the galaxy at the time of bar formation. There is a mild correlation between the bar formation time and the stellar-to-dark matter ratio at $z=0$, while there is no significant correlation between bar formation time and stellar mass at $z=0$ (i.e no obvious signs of mass downsizing within this mass range). There is a trend of decreasing stellar mass for earlier bar formation time (right panel), as expected, since galaxies at higher redshifts have lower stellar masses.
@@
Figure 9: The bar length-to-disc size ratio as a function of the formation time of the bar for the non-interacting galaxies. Left: The ratio at the time of bar formation; Middle: The ratio at $z=0$; Right: The change in the ratio from bar formation time to $z=0$. We see that the ratio of bar length-to-disc size is larger at the time of bar formation for early forming bars, and decreases at $z=0$ due to the increasing disc size.
@@ 2.2 Analysis
For estimating the bar fraction in Fig. 4 we also visually inspect the galaxies, to ensure that when $A_{2}\geq 0.25$, this is not due, for example, to a transient off-centering of the disc due to a merger (which can occur frequently at high redshifts). The aforementioned parameters and cuts used for the bar strength are applied such that our definition of ‘barred galaxy’ agrees well with tests that were done in which the authors visually identified bars in the simulations. We note that slight changes in these parameters do not change the overall conclusions of this work.
@@
Bar length, $R_{\rm bar}$: There are various methods for estimating the length of the bar, from ellipse-fitting methods used widely in observational studies, to methods using the $m=2$ Fourier mode of the surface density, which are more commonly employed in numerical studies (see e.g. Erwin 2005; Gadotti 2011; Hilmi et al. 2020; Fragkoudi et al. 2021; Petersen et al. 2023). Here we use a method which is well-suited for estimating the bar length in an automated way for simulated galaxies, across a large number of snapshots, i.e. we use the radius at which $A_{2}$ falls below 70% of its maximum value. The bar length at the time of bar formation can vary rapidly, in particular when the bar forms due to an interaction or after a merger. Therefore, when estimating the bar length at the time of formation, we take the average over three snapshots after $t_{\rm bf}$. We note that we have tested various methods for obtaining the bar length, such as a definition based on the constancy of the phase of the $m=2$ Fourier mode. We found this latter method to be more prone to noisy bar length estimates, hence why we use the former method here (see also Figure B2 of Fragkoudi et al. 2021).
@@
The galaxy assembly time, $t_{\rm gal,50}$: is defined as the time at which the galaxy assembles 50% of its stellar mass within a radius of 20 kpc.
$M_{\rm\star}$: In what follows, we calculate the stellar mass of the simulated galaxy as the mass within 10% of the virial radius, unless otherwise stated.
@@ 3.1 Global Properties of Barred Galaxies
Mass Build-Up and Bar Formation Time In the top panel of Fig. 5 we show the bar strength at $z=0$ as a function of the bar formation time. We find a weak correlation between the bar strength at $z=0$ and the age (or formation time) of the bar, indicating that older bars tend to be slightly stronger. In the second panel of the figure we show the bar formation time as a function of the time at which the galaxy becomes baryon dominated (see Section 2.2 for details). We see that there is a correlation between these quantities, suggesting that galaxies form a bar when the galaxy passes a threshold of baryon-dominance (and see also Section 4.1).
In the bottom panel of Fig. 5 we show the bar formation time as a function of the formation time of the galaxy (which we define as the time at which the galaxy assembles 50% of its stellar mass). We find a strong correlation between the two parameters, indicating that younger bars tend to form in galaxies that assembled their mass later. We note that the yellow point which is an outlier in this relation (Au25) has an underestimated bar age due to the method we use for obtaining the bar formation time. As discussed in Sec. 2.2, the bar age is taken as the last time $A_{2}$ becomes larger than 0.25. This galaxy has a bar strength that oscillates around the cut-off value of 0.25, starting at around $t_{\rm lb}\sim 3.25$ Gyr. The last time it crosses above 0.25 is at $t_{\rm lb}=0.34$ Gyr – making this the bar age according to our definition. However, there is a clearly formed – albeit weak – bar at earlier times already at $t_{\rm lb}\sim 3.25$ Gyr. The dashed grey line indicates a linear fit to the sample with a slope of 1.6 and intercept of -4.7 Gyr.
@@ 3.2 Evolution of Bar Properties
In Figure 7 we show the evolution of bar length as a function of time. In particular, we group together bars according to their formation time, i.e. we select bars that form between lookback times of $t_{\rm lb}=0-2.7$ Gyr ago (dark purple), $t_{\rm lb}=2.8-4.9$ Gyr (dark blue), $t_{\rm lb}=5-7.1$ Gyr (dark green) and $t_{\rm lb}=7.2-11$ Gyr (light green), as indicated by the legend. These bins are chosen such that they split the sample into 4 groups with approximately the same number of galaxies. We find that bars that form at low redshifts, tend to form shorter bars which subsequently grow longer over time.
On the other hand, bars that form at higher redshifts (e.g. around $z\sim 1$ – see the green line) tend to already be quite long when they form, and their length does not evolve much over the galaxy’s lifetime. We hypothesize that this is due to the fact that galaxies that form bars at high-$z$ are triggered by a merger, and therefore form long and “saturated”, i.e. they are not able to grow by exchanging more angular momentum. Indications of this can be seen in the inset of Fig. 7, where we show the time of one of the three most massive mergers in the galaxy, versus the time of bar formation. We see that the galaxies that form bars at high redshift all do so around the time of one of their most massive mergers. The growth of bars is mediated through the exchange of angular momentum between the inner disc and the outer disc and halo, with the disc acting as an ‘emitter’ and the halo acting as a ‘sink’ of angular momentum (see e.g. Athanassoula 2003). We speculate that if a bar forms through a merger/interaction, the torques induced in this process might reduce the amount of angular momentum able to be exchanged. This could be because of the torques induced in the disc which can trigger the bar to form in the first place (see e.g. Łokas 2018; Merrow et al. 2024) or it could be due to the injection of angular momentum in the halo via the interaction/merger, which therefore means the halo is able to absorb less angular momentum. These processes could therefore lead to the angular momentum exchange being ‘saturated’, which can impede the further growth of the bars. While there are hints of this in the simulations (e.g. the fact that the bars forming at high redshifts all form around the time of a significant merger, and that they do not grow longer in time), this hypothesis would need to be investigated in more detail, which we defer to future work.
@@ 4.1 Baryon dominance
Figure 11: The average (solid lines) and standard deviation (shaded regions) of the ratio of stellar-to-dark matter mass within 5 kpc as a function of time for the barred galaxies (excluding the interacting galaxies), in red, and for the unbarred galaxies, in blue. On average, for barred galaxies, the stellar mass dominates over the dark matter mass within the central 5 kpc by $z\sim 0$, which is in contrast to unbarred galaxies which tend to be more dark matter dominated over all cosmic times.
[…]
@@
In Fig. 11 we explore the ratio of stellar-to-dark matter within 5 kpc ($f_{\rm bd}$) as a function of time; we find that, already at earlier lookback times, the barred galaxies are more baryon dominated than the unbarred sample which remains dark matter dominated even at low redshifts. This trend also holds if we explore the baryon dominance within the half-mass radii of the galaxies ($r_{\rm 50}$), as is shown in Fig. 22, with the differences between the two populations being even more pronounced (due to the fact that the half-mass radii of barred galaxies are smaller than for unbarred galaxies in our sample).
In Fig. 12 we show the growth of the stellar mass of the barred (red) and unbarred (blue) galaxies as a function of time. As discussed in Section 2.2, we calculate the stellar mass within 10% of the virial radius of the galaxy. We find that barred galaxies assemble their stellar mass more rapidly than unbarred galaxies (see also Rosas-Guevara et al. 2020). We quantify this by exploring when the samples assemble 50% of their stellar mass, and find that barred galaxies assemble their stellar mass on average $\sim$1 Gyr earlier than unbarred galaxies. Barred galaxies therefore tend to grow their stellar component more rapidly than unbarred galaxies, forming an early massive stellar disc, which is conducive to bar formation.
@@ 5 Discussion
Figure 17: The stellar Tully-Fisher relation at $z=0$ for galaxies in Auriga barred galaxies (excluding interacting) and unbarred galaxies. Left: The grey dot-dashed line shows the relation given by fitting all galaxies together, while the shaded region indicates the intrinsic vertical scatter $\sigma_{y}$. Right: The red (blue) dashed line shows the relation when fitting only the barred (unbarred) galaxies. The shaded regions indicate the uncertainties from boostrapping (see text).We find that the TF relation for barred galaxies is offset from that of the unbarred galaxies.
Figure 18: The multi-dimensional parameter space of bar formation: dark matter halo mass ($M_{\rm DM}$), stellar mass ($M_{\star}$), baryon-dominance ($f_{\rm bd}$), fraction of accreted stars in the central bulge regions ($f_{\rm acc}$), the Toomre parameter within the half-mass radius ($Q_{r_{50}}$) and the gas fraction ($f_{\rm gas}$), all at $z=0$. The red (blue) points indicate the barred (unbarred) sample, including interacting galaxies. We see that the barred-unbarred populations separate out for the baryon dominance, accreted bulge fraction and Toomre $Q$ parameters.
@@ 5.1 Implications for the Tully-Fisher relation
The Tully-Fisher relation (TF; Tully & Fisher 1977) is an empirical relation between the rotation velocity of galaxies – often the radius at which the rotation curve flattens out – versus their luminosity. This relation has low intrinsic scatter, making it a powerful tool for determining distances to galaxies, while it has also proved useful for testing galaxy formation models in $\Lambda$CDM (Mo et al., 1998) and theories of modified gravity, such as Modified Newtonian Dynamics, or MOND (see e.g. Milgrom 1983; McGaugh 2012). In Fig. 17 we explore the consequences of our findings in Section 4.1 – i.e. the fact that barred galaxies tend to be more baryon-dominated than unbarred galaxies – on the stellar TF relation (STFR; in Appendix Bar formation and evolution in the cosmological context: Inputs from the Auriga simulations we show results for the Baryonic Tully Fisher relation). We obtain the circular velocity at $z=0$ of the galaxies in our sample at 30 kpc – which corresponds roughly to the flat part of the rotation curve for the galaxies we are exploring, i.e. $V_{\rm flat}$. We then fit the relation $\log{M_{\star}}=m\log{V_{\rm flat}}+y_{0}$ to our sample of barred galaxies (excluding interracting galaxies) and unbarred galaxies. In the left panel of Fig. 17 we show a least squares fit to this sample (grey dashed curve). The vertical scatter of the relation in our sample is $\sigma_{y}=0.05$dex, as indicated with the grey shaded region. In the right panel of the Figure we fit separately the barred (red) and unbarred (blue) samples. For these fits, we fix the slope of the relation to be that obtained for the total sample and we perform bootstrapping to assess the uncertainties in the fits. The 95% percentile uncertainty in the intercept, as obtained form our fits, is shown with the shaded regions. We find an offset in the STFR of the barred galaxies as compared to that of the unbarred galaxies, of the order of $\Delta y_{0}=0.05$dex. This offset is comparable to the intrinsic scatter of the entire sample, suggesting that, while challenging, it could in principle be observed.